Determine of the graphs of the 2 equations are perpendicular?
y=7x-9
2y=7-x

Answer by nerdybill(7384)

y=7x-9
2y=7-x
.
First, convert both to "slope-intercept" form of
y = mx + b
.
y=7x-9
y=(-1/2)x + 7/2
.
Slope of equation 1 is 7
Slope of equation 2 is -1/2
.
If the two lines are perpendicular, multiplying the two slopes together should equal -1.
.
7 * (-1/2) = -7/2
.
Since it does not, the two lines are NOT perpendicular.